# Differentially Private Zeroth-Order Methods for Scalable Large Language Model Fine-tuning

### v4

@@ A.1. Complete Proof
(27) $\begin{split}&\mathbb{E}\left[\left\|\nabla f_{\beta}(\boldsymbol{\theta})-\mathbb{E}_{\mathbf{v}}\left[\nabla f_{\beta}(\boldsymbol{\theta})\right]\right\|^{2}\right]\\ \leq&2d\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v})-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}+\beta\mathbf{v})\right]\right)\right\|^{2}\right]\\ &+2d\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}-\beta\mathbf{v})-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}-\beta\mathbf{v})\right]\right)\right\|^{2}\right]\leq 64d\beta^{2}L^{4}.\end{split}$
Following assumption 1 that $\mathbb{E}\left[\left\|\nabla F_{\beta}(\boldsymbol{\theta})-\nabla f_{\beta}(\boldsymbol{\theta},x)\right]\right\|^{2}\leq\gamma^{2}$
(28) $\begin{split}&\left\|\mathbb{E}_{\mathbf{v}}\left[\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})\right]-\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ \leq&\frac{1}{P^{2}m^{2}}\sum_{p=1}^{P}\sum_{i=1}^{m}\left\|\nabla F_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ &+\frac{64d\beta_{s}^{2}L^{4}}{Pm}\\ \leq&\frac{1}{P^{2}m^{2}}\sum_{p=1}^{P}\sum_{i=1}^{m}\left\|\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ &+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}.\\ \end{split}$
@@
(29) $\begin{split}\mathbb{P}\left(\frac{1}{2\beta}\left|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right|\geq C\right)\leq 2e^{-\frac{C^{2}}{2L^{2}}}.\end{split}$
@@
(30) $\begin{split}&\mathbb{E}\left[\left\|\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\right]\\ =&\left\|\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-C\right\|^{2}\mathbb{P}(\overline{Q_{t+1}^{i}})\\ \leq&\left(2dC^{2}+4dL^{2}\right)\cdot\exp\left(-\frac{C^{2}}{2L^{2}}\right).\\ \end{split}$
@@
(31) $\begin{split}&\mathbb{E}\left[\left\|\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\right]\leq\frac{8dC^{2}}{e}.\\ \end{split}$
@@
(32) $\begin{split}\mathbb{E}[||\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})&-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]\\ \leq&\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{8dC^{2}}{ePm}+\frac{\gamma^{2}}{Pm}.\end{split}$
Lemma A.2 (Restatation of Lemma 4.3).
Under Assumption 1 and $f(\boldsymbol{\theta},x)$ is a $\rho$-weakly-convex function of $\boldsymbol{\theta}$, by applying DP-ZOO (Algorithm 2) with $\eta_{s}\leq\frac{\beta_{s}}{\sqrt{d}L}$, for any $\boldsymbol{\theta}\in\mathcal{R}^{d}$, we have
(33) $\begin{split}\mathbb{E}[\phi_{s}(\hat{\boldsymbol{\theta}}_{T_{s}})&-\phi_{s}(\boldsymbol{\theta})]\leq\left(\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{8dC^{2}}{ePm}+\frac{\gamma^{2}}{Pm}\right).\\ \end{split}$
Proof for Lemma 4.3: Recall that $\phi_{s}(\boldsymbol{\theta})=f_{\beta_{s}}(\boldsymbol{\theta})+\frac{1}{2\lambda}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{s}\right\|^{2}$, let $F_{\beta_{s}}(\boldsymbol{\theta})=f_{\beta_{s}}(\boldsymbol{\theta},\mathcal{D})$, $r_{s}(\boldsymbol{\theta})=\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{s}\right\|^{2}$. Due to the $\rho$-weak-convexity of $f(\boldsymbol{\theta})$, the $\frac{1}{\lambda}$-strong-convexity of $r_{s}(\boldsymbol{\theta})$ and the $\frac{\sqrt{d}L}{\beta_{s}}$-smoothness of $f_{\beta_{s}}(\boldsymbol{\theta})$, we have the following three inequality
(34) $\begin{split}&F_{\beta_{s}}(\boldsymbol{\theta})\geq F_{\beta_{s}}(\boldsymbol{\theta}_{t})+\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t}),\boldsymbol{\theta}-\boldsymbol{\theta}_{t}\rangle-\frac{\rho}{2}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}\\ &r_{s}(\boldsymbol{\theta})\geq r_{s}(\boldsymbol{\theta}_{t+1})+\langle\partial r_{s}(\boldsymbol{\theta}_{t+1}),\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\rangle+\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}\\ &F_{\beta_{s}}(\boldsymbol{\theta}_{t})\geq F_{\beta_{s}}(\boldsymbol{\theta}_{t+1})-\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}_{t}\rangle-\frac{\sqrt{d}L}{2\beta_{s}}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}\right\|^{2}.\\ \end{split}$
@@
(35) $\begin{split}&F_{\beta_{s}}(\boldsymbol{\theta}_{t+1})+r_{s}(\boldsymbol{\theta}_{t+1})-(F_{\beta_{s}}(\boldsymbol{\theta})+r_{s}(\boldsymbol{\theta}))\\ \leq&\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})+\partial r_{s}(\boldsymbol{\theta}_{t+1}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}\rangle\\ &+\frac{\rho}{2}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}+\frac{\sqrt{d}L}{2\beta_{s}}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}\right\|^{2}-\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}.\end{split}$
If we set the gradient in $\boldsymbol{\theta}_{t+1}$ to 0, there exists $\partial r_{s}(\boldsymbol{\theta}_{t+1})$ such that
(36) $\partial r_{s}(\boldsymbol{\theta}_{t+1})=\frac{1}{\eta_{s}}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}).$
@@
where $\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})=\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)+\mathbf{z}_{t}^{p}$ and $\mathbf{z}_{t}^{p}=z_{t}^{p}\cdot\mathbf{v}_{t}^{p}\left(z_{t}^{p}\sim N(0,\sigma^{2}C^{2})\right)$. Plugging the above equation into 35 and setting $\hat{\boldsymbol{\theta}}_{t+1}$ be the updated parameter on dataset $\mathcal{D}$ without DP noise in iteration $t+1$. Since $\eta$ is decreasing and $\beta$ is increasing, $\eta_{s}\leq\frac{\beta_{s}}{\sqrt{d}L}$ holds true for all $s\in[0,S]$ when $\eta_{1}\leq\frac{\beta_{1}}{\sqrt{d}L}$, by taking $\eta_{s}\leq\frac{\beta_{s}}{\sqrt{d}L}$, we have
@@
(37) $\begin{split}&F_{\beta_{s}}(\boldsymbol{\theta}_{t+1})+r_{s}(\boldsymbol{\theta}_{t+1})-(F_{\beta_{s}}(\boldsymbol{\theta})+r_{s}(\boldsymbol{\theta}))\\ \leq&\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}\rangle+\langle\frac{1}{\eta_{s}}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}\rangle\\ &+\frac{\rho}{2}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}+\frac{\sqrt{d}L}{2\beta_{s}}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}\right\|^{2}-\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}\\ =&\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}),\boldsymbol{\theta}_{t+1}-\hat{\boldsymbol{\theta}}_{t+1}+\hat{\boldsymbol{\theta}}_{t+1}-\boldsymbol{\theta}\rangle\\ &+\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}\\ =&\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}),\hat{\boldsymbol{\theta}}_{t+1}-\boldsymbol{\theta}\rangle\\ &+\eta_{s}\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})\right\|^{2}+\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}\\ &-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}.\\ \end{split}$
@@
(38) $\begin{split}&\mathbb{E}\left[\phi_{s}(\boldsymbol{\theta}_{t+1})-\phi_{s}(\boldsymbol{\theta})\right]\leq\eta_{s}\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})\right\|^{2}\right]\\ &+\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}.\\ \end{split}$
Following Lemma 4.2 and by taking summation of the above inequality from $t=0$ to $T_{s}-1$, $\lambda<1/\rho$, we have
(39) $\begin{split}&\sum_{t=0}^{T_{s}-1}\phi_{s}(\boldsymbol{\theta}_{t+1})-\phi_{s}(\boldsymbol{\theta})\leq\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}T_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{8dC^{2}}{ePm}+\frac{\gamma^{2}}{Pm}\right)\\ &-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{T_{s}}\right\|^{2}.\end{split}$
By employing Jensens’ inequality, denoting the output of stage $s$ by $\boldsymbol{\theta}^{s}=\hat{\boldsymbol{\theta}}_{T_{s}}=\frac{1}{T_{s}}\sum_{t=1}^{T_{s}}\boldsymbol{\theta}_{t}$, since $\rho\leq 1/\lambda$ we can obtain that
(40) $\begin{split}\phi_{s}(\hat{\boldsymbol{\theta}}_{T_{s}})&-\phi_{s}(\boldsymbol{\theta})\leq\left(\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).\\ \end{split}$
Theorem A.3 (Restatation of Theorem 4.4).
@@
(41) $\phi_{s}(\boldsymbol{\theta}^{S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$
@@
(42) $\mathcal{O}\left(\left(\gamma^{2}+dC^{2}+d\beta_{S}^{2}L^{4}+\frac{dC^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$
Proof of Theorem 4.4.

### v5

@@ A-A Complete Proof
$\begin{split}&\mathbb{E}\left[\left\|\nabla f_{\beta}(\boldsymbol{\theta})-\mathbb{E}_{\mathbf{v}}\left[\nabla f_{\beta}(\boldsymbol{\theta})\right]\right\|^{2}\right]\\ \leq&2d\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v})-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}+\beta\mathbf{v})\right]\right)\right\|^{2}\right]\\ &+2d\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}-\beta\mathbf{v})-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}-\beta\mathbf{v})\right]\right)\right\|^{2}\right]\\ \leq&64d\beta^{2}L^{4}.\end{split}$ (27)
Following assumption 1, we can derive that
$\begin{split}&\small\left\|\mathbb{E}_{\mathbf{v}}\left[\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})\right]-\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ \leq&\frac{1}{P^{2}m^{2}}\sum_{p=1}^{P}\sum_{i=1}^{m}\left\|\nabla F_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ &+\frac{64d\beta_{s}^{2}L^{4}}{Pm}\\ \leq&\small\frac{1}{P^{2}m^{2}}\sum_{p=1}^{P}\sum_{i=1}^{m}\left\|\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ &+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}.\\ \end{split}$ (28)
@@
$\begin{split}\small\mathbb{P}\left(\frac{1}{2\beta}\left|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right|\geq C\right)\leq 2e^{-\frac{C^{2}}{2L^{2}}}.\end{split}$ (29)
@@
$\begin{split}&\mathbb{E}\left[\left\|\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\right]\\ =&\left\|\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-C\right\|^{2}\mathbb{P}(\overline{Q_{t+1}^{i}})\\ \leq&\left(2dC^{2}+4dL^{2}\right)\cdot\exp\left(-\frac{C^{2}}{2L^{2}}\right).\\ \end{split}$ (30)
@@
$\begin{split}&\mathbb{E}\left[\left\|\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})-\mathrm{CLIP}\left(\nabla f_{\beta_{s}}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\right]\leq\frac{8dC^{2}}{e}.\\ \end{split}$ (31)
@@
$\begin{split}\small\mathbb{E}[||\nabla&F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]\\ \leq&\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{8dC^{2}}{ePm}+\frac{\gamma^{2}}{Pm}.\end{split}$ (32)
Lemma 7 (Restatation of Lemma 5).
Under Assumption 1 and $f(\boldsymbol{\theta},x)$ is $\rho$-weakly-convex of $\boldsymbol{\theta}$, by applying DP-ZOO (Algorithm 2) with $\eta_{s}\leq\frac{\beta_{s}}{\sqrt{d}L}$, for any $\boldsymbol{\theta}\in\mathcal{R}^{d}$, we have
$\begin{split}\mathbb{E}&[\phi_{s}(\hat{\boldsymbol{\theta}}_{T_{s}})-\phi_{s}(\boldsymbol{\theta})]\leq\left(\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{8dC^{2}}{ePm}+\frac{\gamma^{2}}{Pm}\right).\\ \end{split}$ (33)
Proof for Lemma 5: Recall that $\phi_{s}(\boldsymbol{\theta})=f_{\beta_{s}}(\boldsymbol{\theta})+\frac{1}{2\lambda}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{s}\right\|^{2}$, let $F_{\beta_{s}}(\boldsymbol{\theta})=f_{\beta_{s}}(\boldsymbol{\theta},\mathcal{D})$, $r_{s}(\boldsymbol{\theta})=\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{s}\right\|^{2}$. Due to the $\rho$-weak-convexity of $f(\boldsymbol{\theta})$, the $\frac{1}{\lambda}$-strong-convexity of $r_{s}(\boldsymbol{\theta})$ and the $\frac{\sqrt{d}L}{\beta_{s}}$-smoothness of $f_{\beta_{s}}(\boldsymbol{\theta})$, we have the following three inequality
$\begin{split}&F_{\beta_{s}}(\boldsymbol{\theta})\geq F_{\beta_{s}}(\boldsymbol{\theta}_{t})+\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t}),\boldsymbol{\theta}-\boldsymbol{\theta}_{t}\rangle-\frac{\rho}{2}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}\\ &r_{s}(\boldsymbol{\theta})\geq r_{s}(\boldsymbol{\theta}_{t+1})+\langle\partial r_{s}(\boldsymbol{\theta}_{t+1}),\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\rangle+\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}\\ &\footnotesize F_{\beta_{s}}(\boldsymbol{\theta}_{t})\geq F_{\beta_{s}}(\boldsymbol{\theta}_{t+1})-\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}_{t}\rangle-\frac{\sqrt{d}L}{2\beta_{s}}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}\right\|^{2}.\\ \end{split}$ (34)
@@
$\begin{split}&F_{\beta_{s}}(\boldsymbol{\theta}_{t+1})+r_{s}(\boldsymbol{\theta}_{t+1})-(F_{\beta_{s}}(\boldsymbol{\theta})+r_{s}(\boldsymbol{\theta}))\\ \leq&\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})+\partial r_{s}(\boldsymbol{\theta}_{t+1}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}\rangle\\ &+\frac{\rho}{2}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}+\frac{\sqrt{d}L}{2\beta_{s}}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}\right\|^{2}-\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}.\end{split}$ (35)
If we set the gradient in $\boldsymbol{\theta}_{t+1}$ to 0, there exists $\partial r_{s}(\boldsymbol{\theta}_{t+1})$ that
$\partial r_{s}(\boldsymbol{\theta}_{t+1})=\frac{1}{\eta_{s}}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}).$ (36)
@@
where $\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})=\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}($$\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},$ $x_{t+1}^{i}$ ))))) ) $+\mathbf{z}_{t}^{p}$ and $\mathbf{z}_{t}^{p}=z_{t}^{p}\cdot\mathbf{v}_{t}^{p}\left(z_{t}^{p}\sim N(0,\sigma^{2}C^{2})\right)$. Plugging the above equation into 35 and setting $\hat{\boldsymbol{\theta}}_{t+1}$ be the updated parameter on dataset $\mathcal{D}$ without DP noise in iteration $t+1$. Since $\eta$ is decreasing and $\beta$ is increasing, $\eta_{s}\leq\frac{\beta_{s}}{\sqrt{d}L}$ holds true for all $s\in[0,S]$ when $\eta_{1}\leq\frac{\beta_{1}}{\sqrt{d}L}$, by taking $\eta_{s}\leq\frac{\beta_{s}}{\sqrt{d}L}$, we have
$F_{\beta_{s}}(\boldsymbol{\theta}_{t+1})+r_{s}(\boldsymbol{\theta}_{t+1})-(F_{\beta_{s}}(\boldsymbol{\theta})+r_{s}(\boldsymbol{\theta}))$ (37)
$\leq$ $\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}\rangle$
$+\langle\frac{1}{\eta_{s}}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}),\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}\rangle+\frac{\rho}{2}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}$
@@
$+\frac{\sqrt{d}L}{2\beta_{s}}\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t+1}\right\|^{2}-\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}$
$=$ $\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}),\boldsymbol{\theta}_{t+1}-\hat{\boldsymbol{\theta}}_{t+1}+\hat{\boldsymbol{\theta}}_{t+1}-\boldsymbol{\theta}\rangle$
$+\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}$
$\leq$ $\langle\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1}),\hat{\boldsymbol{\theta}}_{t+1}-\boldsymbol{\theta}\rangle$
$+\eta_{s}\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})\right\|^{2}$
$+\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}.$
@@
$\begin{split}&\small\mathbb{E}\left[\phi_{s}(\boldsymbol{\theta}_{t+1})-\phi_{s}(\boldsymbol{\theta})\right]\leq\eta_{s}\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})\right\|^{2}\right]\\ &\small+\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}\right\|^{2}-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{t+1}\right\|^{2}.\\ \end{split}$ (38)
Following Lemma 4 and by taking summation of the above inequality from $t=0$ to $T_{s}-1$, $\lambda<1/\rho$, we have
$\begin{split}&\sum_{t=0}^{T_{s}-1}\phi_{s}(\boldsymbol{\theta}_{t+1})-\phi_{s}(\boldsymbol{\theta})\leq\left(\frac{1}{2\eta_{s}}+\frac{\rho}{2}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}T_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{8dC^{2}}{ePm}+\frac{\gamma^{2}}{Pm}\right)\\ &-\left(\frac{1}{2\eta_{s}}+\frac{1}{2\lambda}\right)\left\|\boldsymbol{\theta}-\boldsymbol{\theta}_{T_{s}}\right\|^{2}.\end{split}$ (39)
By employing Jensens’ inequality, denoting the output of stage $s$ by $\boldsymbol{\theta}^{s}=\hat{\boldsymbol{\theta}}_{T_{s}}=\frac{1}{T_{s}}\sum_{t=1}^{T_{s}}\boldsymbol{\theta}_{t}$, we can obtain that
$\begin{split}&\phi_{s}(\hat{\boldsymbol{\theta}}_{T_{s}})-\phi_{s}(\boldsymbol{\theta})\leq\left(\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).\\ \end{split}$ (40)
Theorem 5 (Restatation of Theorem 2).
@@
$\phi_{s}(\boldsymbol{\theta}^{S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$ (41)
@@
$\mathcal{O}\left(\left(\gamma^{2}+dC^{2}+d\beta_{S}^{2}L^{4}+\frac{dC^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$ (42)
Proof of Theorem 2.